OpenStudy (katwqyz):

Solve the system of equations by substitution. x = -y + 5 3x + y = 25

OpenStudy (3mar):

May I help?

OpenStudy (daniellelovee):



likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\) what do u think it is ?

OpenStudy (katwqyz):

dunno @likeabossssssss

RhondaSommer (rhondasommer):

\[ x = -y + 5\\3x + y = 25\] \[3(-y+5)+y=25\\-3y+15+y=25\\-2y+15=25\\-2y=10\\y=-5\] \[x=-(-5)+5\\x=5+5\\x=10\] \[x=10\\y=-5\\(x,y)\\(10,-5)\]

OpenStudy (katwqyz):

thanks so much <3 <3 <3